### v2

# Toward Signed Distance Function based Metamaterial Design: Neural Operator Transformer for Forward Prediction and Diffusion Model for Inverse Design

Abstract
The inverse design of metamaterial architectures presents a significant challenge, particularly for nonlinear mechanical properties involving large deformations, buckling, contact, and plasticity. Traditional methods, such as gradient-based optimization, and recent generative deep-learning approaches often rely on binary pixel-based representations, which introduce jagged edges that hinder finite element (FE) simulations and 3D printing. To overcome these challenges, we propose an inverse design framework that utilizes a signed distance function (SDF) representation combined with a conditional diffusion model. The SDF provides a smooth boundary representation, eliminating the need for post-processing and ensuring compatibility with FE simulations and manufacturing methods. A classifier-free guided diffusion model is trained to generate SDFs conditioned on target macroscopic stress-strain curves, enabling efficient one-shot design synthesis. To assess the mechanical response and the quality of the generated designs, we introduce a forward prediction model based on Neural Operator Transformers (NOT), which accurately predicts homogenized stress-strain curves and local solution fields for arbitrary geometries with irregular query meshes. This approach enables a closed-loop process for general metamaterial design, offering a pathway for the development of advanced functional materials.
Keywords: Diffusion model; Inverse design; Neural operator transformer; Arbitrary geometry; Metamaterial
1 Introduction
The evolution of material science has been revolutionized by the advent of additive manufacturing, facilitating the synthesis of materials with unprecedented customized properties. Engineers and designers are no longer confined to the limited options presented by natural materials but now have the ability to explore the vast design and property spaces offered by metamaterials. These engineered materials are meticulously crafted to achieve mechanical properties that were once considered beyond reach. Commonly realized through periodic arrangements of small-scale structural unit cells, metamaterials provide an innovative platform for exploring novel mechanical behaviors. In recent years, the production of artificial materials with bespoke characteristics has garnered significant interest across various engineering disciplines, e.g. bio-inspired material [1], soft robotics [2, 3], nanophotonics [4, 5], automotive [6] and aerospace industries [7]. The flexibility offered by additive manufacturing processes allows for the creation of architected materials, or metamaterials, at diverse scales and sizes. The exceptional mechanical properties of these materials, such as their ability to bend, stretch, compress, or respond to forces, are intricately linked to their internal structure, typically composed of repeating geometric patterns made from conventional materials.
The material constitutive laws that govern the mechanical behavior of metamaterials are well developed and various numerical methods have been used to predict the material properties based on the unit cell structure. However, the inverse design of the unit cell structure that can achieve a desired material property remains an open challenge, especially for nonlinear mechanical properties involving large deformations, structural buckling, frictional contact, and plasticity. Traditional methods rely on a trial-and-error approach or gradient-based topology optimization [8, 9], and genetic algorithms [10]. These approaches are computationally expensive, as they often involve iterative solutions of PDEs using numerical methods like finite element methods (FEM). Moreover, the efficiency of search algorithms declines significantly as the design space grows. Although such computational cost for repeatedly solving PDEs can be alleviated by using reduced-order models [11] or surrogate models [12, 13, 14, 15, 16, 17], the gradient-based optimization process may be trapped in local minima [18].
To overcome these challenges, numerous probabilistic generative models have been proposed for “one-shot” design, eliminating the need for multiple iterations. These include conditional generative adversarial networks (cGAN) [19, 20], variational autoencoders (VAE) [21], conditional variational autoencoders (cVAE) [22], univariate conditional variational autoencoder (UcVAE) [18], and conditional diffusion models [23, 24, 25, 26]. By learning the underlying data distribution, these models generate diverse solutions within the design space based on a given design target.
However, all the aforementioned methods, whether gradient-based optimization or generative deep-learning models, rely on a binary pixel-based representation of metamaterial architectures. This representation naturally introduces jagged edges (as shown in Fig. S1(a)), which pose challenges when applied directly in finite element (FE) simulations, potentially causing significant mesh distortions that affect both convergence and accuracy [26]. Additionally, the jagged boundaries complicate fabrication via 3D printing, as sharp edges can lead to structural defects and material failure. To address these issues, a boundary-smoothing process is typically applied to the designed binary pixel-based geometry [23, 26, 27]. Furthermore, accurately representing geometry with a binary pixel-based approach requires high-resolution grids, significantly increasing computational costs for design.
Instead of relying on a binary pixel-based approach, we propose an inverse design framework for metamaterial architectures based on a signed distance function (SDF) and a conditional diffusion model. Specifically, a SDF maps a coordinate x to a scalar $\phi(\textbf{x})$, which represents the signed distance to the nearest point on the geometry’s surface. A negative value indicates a point outside the geometry, while a positive value denotes a point inside. Consequently, the level set $\phi(\textbf{x})=0$ defines the shape’s boundary (Fig. S1(b)). The SDF representation is much more accurate than the binary pixel-based representation. This smooth boundary representation eliminates the need for additional boundary smoothing, making it directly applicable to finite element (FE) simulations and 3D printing. While the SDF representation has been utilized in level-set-based topology optimization methods [28, 29, 30], these approaches are computationally expensive and face challenges in generating new holes or complex topologies—limitations that are effectively addressed by generative AI methods. To design the SDF, we employ a classifier-free guided diffusion model, conditioning the design process on a given target stress-strain curve. The diffusion model learns to synthesize the SDF from pure Gaussian noise, and the resulting SDF is then used to extract the geometry using a marching algorithm. Additionally, our method can design more general unit cell geometries by incorporating Gaussian random fields with the periodic Fourier method to generate the geometry data. This approach contrasts with previous works [23, 25], which generate one-quarter of the unit cell and mirror it twice to obtain the full unit cell, thus limiting the designs to bi-axis symmetrical geometries.
As generative models are inherently stochastic, they can generate multiple solutions for a single target. This one-to-many design capability is particularly useful for exploring the design space and identifying optimal solutions. However, it is impractical to run FE simulations for each generated geometry to evaluate their mechanical properties and material response. Instead, a forward surrogate model is a prevalent choice to replace the repeated FE simulations.
Neural operators have shown promising results in mapping infinite-dimensional input functions (e.g., arbitrary geometries) to output functions on query points (e.g., stress-strain curves, solution fields). Notably, the Fourier Neural Operator (FNO) [31] and the Deep Operator Network (DeepONet) [32] have been effective in this domain. The Fourier Neural Operator is designed for fixed, regular meshes and requires a full mesh graph as input. While DeepONet can predict the solution field at arbitrary query points, training DeepONet requires that the number and location of query points remain fixed across samples. For arbitrary geometries, the query points typically come from irregular meshes, and their number varies across samples. Training DeepONet in such cases is challenging and often requires resampling the query points to a fixed number, as demonstrated in Geom-DeepONet [33]. However, Geom-DeepONet is restricted to parametric geometries (e.g., length, thickness, radius, etc.) and does not generalize to fully arbitrary shapes. Moreover, it requires both the geometric parameters (e.g., length, thickness, radius, etc.) and the SDF as inputs. Although recent works like Geo-FNO [34] and geometry-informed FNO (GI-FNO) [35] extend to arbitrary geometries, Geo-FNO still requires a fixed number of query points during training. Moreover, both GI-FNO and Geo-FNO rely on projecting an irregular mesh onto a fixed regular mesh and then mapping it back. Subsequently, researchers have utilized the attention mechanism of transformers [36] in neural operators to effectively fuse information from input functions and query points [37, 38]. Neural Operator Transformers (NOT) are particularly well-suited for irregular meshes due to the attention mechanism and allow for the prediction of solution fields at arbitrary query points. Unlike DeepONet, which uses a simple dot product to fuse information from input functions and query points, the attention mechanism in NOT guides each query point to focus on the relevant information from the input functions. This allows NOT to generalize and capture more complex relationships between input functions and query points.
In this work, we propose forward NOT models to predict the macroscopic stress-strain curve and local solution fields for arbitrary geometries generated by the inverse diffusion model, given a target stress-strain curve. For stress-strain curve prediction, the query points correspond to fixed strain steps, whereas for solution field prediction, the query points consist of irregular mesh nodes. To handle varying numbers and locations of query points across different geometries, NOT is trained on batches of geometries by padding the query points to the maximum count within each batch, analogous to managing variable sequence lengths in natural language processing (NLP).
This manuscript summarizes the key concepts, methodologies, and results of the closed-loop process for inverse metamaterial design and forward prediction.
2 Results
2.1 Data generation
Figure 1: Workflow for data generation. Initially, a 2D periodic Gaussian random field is generated. The Marching algorithm is then used to extract the contour at a random threshold value, defining the boundary of the 2D periodic unit cell, with field values below the threshold set as void. Periodic boundary conditions and compression are applied to the unit cell for Abaqus simulation to obtain the macroscopic stress-strain curve. To augment the data, the unit cell is randomly shifted multiple times, creating different geometries with the same stress-strain curve. The corresponding stress and displacement solution fields are also shifted for the new geometries, allowing for data augmentation without additional FE simulations.
The data-driven machine learning approach requires a large dataset of paired training data. Fig. 1 provides an overview of the data generation process adopted in this work. To generate a large dataset of paired geometry, macroscopic stress-strain curves, and local solution fields, we first generate a 2D periodic Gaussian random field (GRF) with $64\times 64$ pixels on a square domain, incorporating the periodic Fourier method. The GRF is then thresholded at a random value to extract the contour using a marching algorithm, where the extracted contour defines the unit cell boundary of the metamaterial. The GRF values less than the threshold are set as void. Connecting the contours from the marching algorithm with the boundary edges of the square whose GRF values are higher than the threshold provides a closed manifold geometry. We also evaluate the signed distance function (SDF) of the geometry on a regular and uniform $120\times 120$ grid, which represents the geometry as input or output of the deep learning models. This method generates arbitrary, random, more general periodic unit cell geometries and a vast design space.
The generated geometries are then imported into the commercial finite element code ABAQUS® 2024 to compute the macroscopic stress-strain curve and the local solution fields. We fix the bottom of the unit cell along the y-direction ($u_{y}=0$) and deform the unit cell by applying a vertical displacement to the top edge of the unit cell with a maximum applied compressive strain of 0.2. Periodic boundary conditions are applied along the left and right sides of the unit cell. The horizontal displacement of the node nearest to the center of the unit cell is set at $u_{x}=0$ to remove the rigid body motion. Self-contact is enabled to prevent the unit cell from collapsing during compression. Simulation details are provided in the methods Section 4.2.
Figure 2: Overview of model architectures. The inverse denoising diffusion model (b) is used to design the micro-structure geometry represented by SDF, given the target stress-strain curve. The noise estimator is a residual U-Net architecture, and each residual block takes as input the image from previous layer, the time step $t$, and the stress-strain curve. The iterative denoising process generates the SDF from pure Gaussian noise, from which the geometry is extracted using a marching algorithm. Two forward neural operator transformers (a) are developed to predict the macroscopic stress-strain curve and local solution fields for arbitrary geometries. A residual U-net encodes the geometry represented by SDF into key (K) and value (V) for the attention mechanism. The query points, either the strain $\varepsilon$ or node coordinates $(x,y)$, are encoded using NeRF positional encoding [39] and a multi-layer perceptron (MLP) to form the query (Q) for the attention mechanism. This mechanism fuses the geometry information and query points information, guiding each query point to focus on relevant geometry information. The output of a few attention blocks is then decoded using an MLP to the solution fields on the query points.
Based on this method, we randomly generate 10k geometries and their corresponding stress-strain curves and solution fields (e.g., Mises stress, displacement). Extended Data Fig. 1 illustrates the generated dataset, where (a) shows the volume fraction of the 10k samples, (b) displays 64 examples of the generated geometries, and (c) presents their corresponding stress-strain curves.
To alleviate overfitting, we augment the dataset by randomly shifting the unit cell in the x direction multiple times, as shown in Fig. 1. This shifted geometry exhibits the same material response as the original geometry due to the periodic nature of the geometry and boundary conditions. The solution fields of the shifted geometry can be derived by shifting the original solution fields accordingly. With this data augmentation, we have about 73k samples for training and testing the deep learning models.
2.2 Forward prediction
Figure 3: Comparison between the true and predicted stress-strain curves of the test data, arranged from best (left) to 99th percentile (right). The first row shows the geometries while the second row presents the corresponding macroscopic stress-strain curves.
Figure 4: Comparison of the Mises stress and displacement predictions with FE ground truth for the median case of the test data at different strain $\varepsilon$ steps. The first row shows the true Mises stress under the true deformed shape, the second row shows the predicted Mises stress under the predicted deformed shape, and the third row shows the absolute error of the Mises stress under the true deformed shape.
Neural operator transformers (NOT) map infinite-dimensional input functions to output functions on arbitrary query points:
$G_{\theta}:\mathcal{F}\rightarrow\mathcal{G}(\textbf{x}),$ (1)
where ${G}_{\theta}$ is the neural operator with learnable parameter $\theta$, $\mathcal{F}$ is the space of input functions (e.g., arbitrary geometries), and $\mathcal{G}$ is the space of output functions (e.g., stress-strain curves, solution fields) on query points x.
The attention mechanism [36] in transformers allows the model to fuse information from the input functions and the query points, guiding each query point to focus on the relevant information from the input functions. It can be described as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{e}}}\right)V,$ (2)
where $Q\in\textbf{R}^{n\times d_{e}}$ is the Query matrix, and $K,V\in\textbf{R}^{n_{k}\times d_{e}}$ are the Key and Value matrices, $d_{e}$ is the embedding dimension, $n$ is the number of query points, and $n_{k}$ is the length of the key and value.
We developed two neural operator transformers for forward prediction solutions from the unit cell geometry represented by SDF, as shown in Fig. 2(a). The first model predicts the stress-strain curve, where the query points are the strain steps. The second model predicts the solution fields (e.g., Mises stress, displacement) at 26 strain steps, with query points being the irregular mesh nodes. Since the SDF is defined on a regular and uniform grid, we use a residual U-Net architecture to encode the geometry into key (K) and value (V) for the attention mechanism. The query points, either the strain $\varepsilon$ or node coordinates $(x,y)$, are encoded using NeRF (Neural Radiance Fields) positional encoding [39] and MLP to the Query (Q) for the attention mechanism. To stabilize training, residual connections and layer normalization are applied in each attention block. The output of a few attention blocks is decoded to the solution fields on the query points. We provide details on the methods and implementations in the Method Section 4 and the supplementary Section S3.
The performance of the forward model for predicting the stress-strain curve from arbitrary unit cell geometries is illustrated in Fig. 3, which shows a comparison between the true and predicted stress-strain curves of the test data, arranged in terms of $L_{2}$ relative error from best (left) to the 99th percentile (right). Further performance details for predicting the stress-strain curve are presented in Fig. S4, which shows the MSE loss of training history and the $L_{2}$ relative error distribution of the test data. The overall mean $L_{2}$ relative error of the test data is 2.6% with a standard deviation of 2.4%, indicating the model’s high accuracy in predicting the stress-strain curve.
We also demonstrate in Fig. 4 the performance of the forward model in predicting the Mises stress and displacement fields for the median case of the test data. The first row displays the true (FE-predicted) Mises stress field plotted on the true deformed shape at different strain steps. The second row shows the predicted (by the NOT) Mises stress under the predicted deformed shape, while the third row presents the corresponding absolute error in Mises stress plotted on the true deformed shape of the unit cell. The $L_{2}$ relative error for the median case is 10.1%. Additionally, we present the best and worst cases of the test data in Fig. S6 and Fig. S7, respectively. The MSE loss of the training history and the $L_{2}$ relative error distribution of the test data for Mises stress and displacement solutions are shown in Fig. S5. The overall mean $L_{2}$ relative error of the test data is 10.3% with a standard deviation of 4.6%, indicating the model’s high accuracy in predicting the solution fields for arbitrary geometries.
2.3 Inverse design by diffusion model
In this section, we describe the classifier-free guided diffusion model [40] trained to design the micro-structure of the unit cell of the metamaterial that can achieve a target stress-strain curve. The label data $y$ is the stress-strain curve obtained from the FE simulation of the corresponding geometry, represented by SDF ($x_{0}$).
Given a sample $\mathbf{x}_{0}$ from a prior data distribution $\mathbf{x}_{0}\sim q(\mathbf{x})$, the forward diffusion process of the denoising diffusion probabilistic model adds a small amount of Gaussian noise to the sample over $T$ steps. This results in a sequence of samples $\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$. Each step in the diffusion process is controlled by a Gaussian distribution,
$q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},~{}\beta_{t}\mathbf{I}\right),\left\{\beta_{t}\in(0,1)\right\}_{t=1}^{T}.$ (3)
This diffusion process has the elegant property that we can sample $\mathbf{x}_{t}$ at any time step $t$ using the reparametrization trick,
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (4)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$.
The reverse diffusion process $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be approximated by a posterior distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\sim\mathcal{N}\left(\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)\right)$, where the variance is $\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{I}$, and $\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)$ is the predicted mean defined as
$\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)\right),$ (5)
in which $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ is the noise estimator approximated by a neural network (NN) with learnable parameters $\theta$.
This reverse generative process is random and not controlled by any specific target. In specific design tasks, we aim to generate fields given the condition of a target such as the stress-strain curve, which requires training the NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ with conditional information. To incorporate the condition information $\mathbf{y}$ into the diffusion process, we use the classifier-free guidance method [40]. In the classifier-free guided diffusion model, the unconditional noise estimators $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\emptyset\right)$ and the conditional one $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ of $p_{\theta}(\mathbf{x}|\mathbf{c})$ are trained in a single NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)$, in which condition information $\mathbf{c}$ is randomly set as $\mathbf{c}=\emptyset$ or $\mathbf{c}=\mathbf{y}$. During the reverse inference process, $\epsilon_{\theta}$ in Eq. 5 is replaced by the linear summation of conditional and unconditional noise estimators,
$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)=(1+w)\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)-w\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\emptyset\right),$ (6)
where $w\geq 0$ is the guidance weight. Further details of the classifier-free guided diffusion model are summarized in the supplementary Section S4.1.
As shown in Fig. 2(b), the noise estimator is approximated by a residual U-Net which takes the input of the current image $x_{t}$, the time step $t$, and the condition information $\mathbf{c}=y$ or $\mathbf{c}=\emptyset$. We use a 2D convolutional layer and an MLP to extract the features of the image from previous level of U-Net and the target stress-strain curve $y$. The time step $t$ is encoded using a time embedding layer and an MLP. The outputs of these three branches are concatenated and fed into a 2D convolutional layer, optionally followed by a self-attention layer. The output is then processed with down-sampling or up-sampling layers for the next level of the U-Net. We provide the detailed methods and implementation in the Method Section 4 and Supplementary Section S4.
Such a generative diffusion model can perform one-to-many designs, where giving a single target stress-strain curve generates multiple unit cell geometries that satisfy the target stress-strain curve. We demonstrate the trained diffusion model for such one-to-many material architecture designs by randomly selecting two target stress-strain curves from the test dataset and feeding them into the inverse diffusion model for 200 geometry solutions. The generated geometries are then fed into the forward model to predict the stress-strain curve. We select the four best design results in terms of $L_{2}$ relative error between the target and prediction of the forward model, as shown in Extended Data Fig. 2(c-f). The comparison between the target and prediction of the forward model is shown in Extended Data Fig. 2(a), with the $L_{2}$ relative error in the legend bracket. The corresponding stress-strain curves obtained using Abaqus simulation are shown in Extended Data Fig. 2(b). The results show that the designed geometries can achieve the target stress-strain curve with high accuracy. With the designed geometry, we can further predict the solution fields (e.g., Mises stress, displacement) at different strain steps using our second forward NOT. The comparison of solution fields for the generated geometry of Extended Data Fig. 2(A)(f) between the true and predicted results is shown in Extended Data Fig. 3.
Figure 5: On-demand one-to-many unit cell design. Two targets of stress-strain curves are generated using the Ramberg-Osgood equation [41]. The top case has material properties of Young’s modulus $E=800$ MPa and reference yield stress $\sigma_{0}=30$ MPa. The bottom case is obtained using $E=1000$ MPa and $\sigma_{0}=40$ MPa, but stress only follows the Ramberg-Osgood equation for $\varepsilon\leq 0.156$. For $\varepsilon\in[0.156,0.2]$, the stress increases linearly with Young’s modulus of $500$ MPa. These two target curves are fed into the inverse diffusion model for 500 SDF solutions with guidance weight $w=1$, which are then fed into the forward models for prediction. We select the four design results with the lowest $L_{2}$ relative error, shown in (c-f). (a) shows the comparison between the target and prediction of (c-f) using the forward model, with the $L_{2}$ relative error indicated in the curve legend. (b) shows the corresponding stress-strain curves obtained using Abaqus.
Figure 6: Comparison of the Mises stress and displacement field predictions with FE ground truth at different strain $\varepsilon$ steps for the designed structure in Fig. 5(A)(e). The first row shows the true Mises stress on the true deformed shape, the second row shows the predicted Mises stress on the predicted deformed shape, and the third row shows the absolute error in Mises stress on the true deformed shape.
While our developed diffusion model performs well on the test dataset for inverse retrieval, real-world applications often demand on-the-fly inverse design of custom-defined material properties with high fidelity. Therefore, instead of using stress-strain curves from FE simulations as shown in Extended Data Fig. 2, a robust inverse design model should be able to generate possible designs from custom and predefined stress-strain curves. Thus, we now use the Ramberg-Osgood equation [41] to define the target stress-strain curve. The Ramberg-Osgood equation is defined as
$\varepsilon=\frac{\sigma}{E}+\alpha\left(\frac{\sigma}{\sigma_{0}}\right)^{n},$ (7)
where $E$ is the Young’s modulus, $\sigma_{0}$ is the yield stress, $\alpha$ is the Ramberg-Osgood coefficient, and $n$ is the hardening exponent. To mimic the self-contact behavior of the material, we modify the Ramberg-Osgood equation to include a critical strain $\varepsilon_{c}$ below which the material follows the Ramberg-Osgood equation and above which the material is assumed to behave as a linear elastic material. The modified Ramberg-Osgood equation is defined as
$\varepsilon=\left\{\begin{array}[]{ll}\frac{\sigma}{E}+\alpha\left(\frac{\sigma}{\sigma_{0}}\right)^{n}&\text{if}~{}\varepsilon\leq\varepsilon_{c},\\ \frac{\sigma-\sigma_{c}}{E^{\prime}}+\varepsilon_{c}&\text{if}~{}\varepsilon>\varepsilon_{c},\end{array}\right.$ (8)
where $\sigma_{c}$ is the stress corresponding to $\varepsilon_{c}$ and $E^{\prime}$ is the Young’s modulus after self-contacting.
Using the Ramberg-Osgood equations, we generate two target stress-strain curves with different material properties, as shown in Fig. 5(A) and (B). Both targets use the same hardening exponent $n=10$ and coefficient $\alpha=0.002$, but differ in Young’s modulus and reference yield stress. Case (A) is obtained using Eq. 7 with a Young’s modulus $E=800$ MPa and yield stress $\sigma_{0}=30$ MPa, while Case (B) is derived from Eq. 8 with $E=1000$ MPa, $\sigma_{0}=40$ MPa, $E^{\prime}=500$ MPa, and $\varepsilon_{c}=0.156$. These target curves are fed into the inverse diffusion model to generate 500 geometry solutions, which are then evaluated using the forward model to predict the stress-strain curves. The four best designs based on the $L_{2}$ relative error are shown in Fig. 5 (c-f). Fig. 5(a) and (b) compare the target and designed stress-strain curves for the four geometries using the forward model and FE simulations, respectively, with the $L_{2}$ relative error indicated in the curve legend. As shown in Fig. 5(a), the designed geometries closely match the custom-defined stress-strain curves. The FE simulations of the designed geometries, shown in Fig. 5(b), confirm the accuracy of the designs. Additionally, we predict the solution fields for the designed geometry in Fig. 5(A)(e) at different strain steps, with results shown in Fig. 6. For case (B), the solution fields for the designed geometry (e) are shown in Extended Data Fig. 5. These results demonstrate that the designed geometries achieve the target stress-strain curves and accurately predict the solution fields.
3 Discussion
The integration of additive manufacturing and advanced computational techniques has opened new avenues for the design and synthesis of metamaterials with tailored mechanical properties with applications in various engineering disciplines including aerospace, automotive, and biomedical. One of the primary challenges addressed in this work is the complexity of inverse design in achieving desired macroscopic mechanical properties, especially in nonlinear scenarios. Traditional methods, such as topology optimization, although effective, are often computationally intensive and prone to local minima due to their reliance on iterative gradient-based optimization schemes. Recently, generative deep learning models have shown potential in inverse design by offering a more efficient approach that eliminates the need for multiple iterations, thus providing a pathway for exploring the vast design space of metamaterials. However, existing methods typically rely on binary pixel-based geometries, which have jagged boundaries and require additional boundary smoothing, and are thus not directly applicable to FE simulations and 3D printing. In contrast, our proposed method designs architected metamaterials using a classifier-free guided diffusion model, representing geometries through signed distance functions (SDFs). This approach enables the generation of multiple SDF solutions for a given target stress-strain curve. The transition from binary pixel-based representations to an SDF framework marks a significant advancement, as SDFs provide smooth boundary representations that eliminate jagged edges. This refinement reduces mesh distortions in finite element simulations and simplifies the 3D printing process, ultimately improving simulation accuracy, convergence, and manufacturability, and ensuring the robustness and reliability of the designed structures.
Real-world applications often require on-the-fly design adaptations to achieve custom material properties. Our results demonstrate the ability of the proposed framework to generate diverse and feasible designs that closely align with custom target stress-strain curves. The stochastic nature of generative diffusion models allows for multiple solutions to a single target, but efficiently evaluating the performance and mechanical behavior of these designs remains a challenge. To address this issue, the forward Neural Operator Transformer (NOT) developed in this work enhances the inverse design framework by enabling accurate prediction of the macroscopic stress-strain curves and the local stress and displacement fields for arbitrary designed geometries. This capability provides deeper insight into the mechanical properties of the designed structures. Additionally, NOT’s adaptability to varying mesh configurations strengthens its applicability, making it a robust prediction tool capable of accommodating the irregularities inherent in real-world applications.
The proposed closed-loop framework, which integrates inverse design and forward prediction models, offers a versatile and efficient approach for designing and synthesizing metamaterials with tailored mechanical properties. Beyond predicting homogenized stress-strain curves, the developed forward neural operator transformer (NOT) models can efficiently compute solution fields for arbitrary geometries. This framework can be naturally extended by conditioning the model not only on target stress-strain curves but also on additional material properties, such as maximum stress, which is crucial for assessing material damage. Expanding the framework primarily requires additional computational efforts, including data generation and model training, while the developed models remain directly applicable.
4 Methods
This section provides details on the generation of random unit cell geometries, FE simulations, and the implementation of deep learning models. Additional explanations can be found in the Supplementary Information.
4.1 Unit cell geometries
To generate periodic unit cells, Bastek et al. [23] generate one-quarter of the unit cell represented by binary pixels and then mirror the other three quarters, which is a common practice in the literature but results in bi-axis symmetric and not fully arbitrary unit cells. Additionally, such binary pixel-based geometries have unsmooth boundaries (Fig. S1(a)), which cannot be directly applied for simulation or 3D printing without a boundary smoothing process. Instead, we generate a random metamaterial unit cell by sampling a periodic 2D Gaussian random field $U\left(x\right)$ on a square domain $[0,1~{}\text{mm}]^{2}$ with a resolution of $64\times 64$, incorporating the periodic Fourier method [42]:
$U\left(x\right)=\sum_{i=1}^{N}\sqrt{2S(k_{i})\Delta k}\left(Z_{1,i}\cdot\cos\left(\textbf{k}_{i}\cdot\textbf{x}\right)+Z_{2,i}\cdot\sin\left(\textbf{ k}_{i}\cdot\textbf{x}\right)\right),$ (9)
where $S$ is the spectrum of the Gaussian covariance model, $Z_{1,i},Z_{2,i}\sim N(0,1)$ are mutually independent and drawn from a standard normal distribution, and $\textbf{k}_{i}$ is the equidistant Fourier grid which ensures the periodic field. This algorithm is implemented in GSTools [43], a Python-based open-source package. The generated Gaussian random field (GRF) is then thresholded at a random value to extract the contour using a marching algorithm, with the extracted contour defining the unit cell boundary of the metamaterial. The GRF values less than the threshold are set as void and the extracted closed contours form the internal boundaries of holes. Unclosed contours intersect with the boundary of the square domain and the segment between the two intersecting points on the same edge of the square domain are connected as additional unclosed boundary contours if the GRF values between the two points are higher than the threshold. Connecting the unclosed contours forms the external boundary of the unit cell. We exclude the generated unit cells containing “islands” and also ensure that opposite boundaries of the domain have enough distance to connect with neighboring unit cells by excluding samples with small edge lengths (less than 8% of domain size). Small internal holes are removed to ensure the geometry is printable. To ensure high mesh quality in Abaqus simulations, adjacent points on a contour are merged if the distance is too close. Using this geometry generation method, we generate about 10,000 valid arbitrary random periodic unit cell geometries. We evaluate the shortest distance of $120\times 120$ uniform grid points in $[-0.1~{}\text{mm},1.1~{}\text{mm}]^{2}$ to the boundary contours for the SDF representation of the geometry.
4.2 FE simulation
We use the commercial finite elements code Abaqus 2024 to simulate the material response of the generated unit cell geometries under compression. The nodes along the bottom edge are fixed along the y-direction ($u_{y}=0$) and the nodes along the top edge are subjected to displacement control with a maximum compressive strain of 0.2. Periodic boundary conditions are applied along the left and right sides of the unit cell. The horizontal motion $u_{x}$ of the nearest node to the center of the unit cell is set at zero to remove rigid body motion. Similar to ref.[23], we apply self-contact with a friction coefficient of 0.4 to prevent the unit cell from collapsing during compression.
An elasto-plastic material model with large deformation is used for the simulations, with material properties detailed in Supplementary Section S2.2. An implicit dynamic solver with virtual mass density $\rho=10^{-8}$ is used to mimic quasi-static compression while allowing better convergence for large deformation and elasto-plastic models [23]. The compression strain is applied progressively in 51 steps. Mixed 1st-order quadrilateral elements and 2nd-order triangular elements with full integration and plane-strain assumption are used for the simulations. The global mesh size is set to 0.04 mm. Note that the mesh size around the boundary contour is approximately 1/64 mm, which is close to the segment length of boundary contours, as the geometry contour is extracted from a GRF with a resolution of $64\times 64$. We performed a mesh sensitivity analysis and found that a mesh size of 0.04 mm is sufficient to capture the material response of the unit cell, as presented in Supplementary Section S2.3. We first use Abaqus CAE with Python API to generate the input files for the Abaqus solver for all geometries and then use the Abaqus solver to perform the simulations based on these input files. All computational tasks of data generation are performed on the DELTA machine at the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign.
4.3 Deep learning models
In this section, we provide a high-level summary of the model architecture. Further implementation details can be found in the supplementary information and our GitHub repository. We developed two forward neural operator transformers (NOT) for predicting the stress-strain curve and the solution fields for arbitrary geometries represented by SDF. The SDF is represented by a 2D image with a shape of $120\times 120$. A 2D U-Net architecture is used to encode the SDF. The U-Net maintains the image size of $120\times 120$ as output, followed by three down-sampling layers to reduce the image size. After reshaping, the output of the geometry encoder is a matrix with a shape of $15^{2}\times d_{e}$, where $d_{e}$ is the embedding dimension.
The other inputs to NOT are the query points, which are the strain steps with a shape of $N\times 1$ or the node coordinates with a shape of $N\times 2$ for the two NOT models, respectively, in which $N$ is the number of query points. For the strain steps, $N=51$. Since the geometry is arbitrary, the number of mesh nodes varies for each geometry. To allow for batch processing, the mesh nodes are padded to the maximum number of nodes in the batch. The NeRF positional encoding followed by an MLP is applied to the query points, whose output has a shape of $N\times d_{e}$, which is then fused with the output of the geometry encoder using the attention mechanism. Residual connections and layer normalization are applied in each attention block to stabilize the training. The output of a few attention blocks is decoded using an MLP to the solution on the query points, e.g., the effective stress with a shape of $N\times 1$ and the 26 frames of solution fields (Mises stress and two displacements) with a shape of $N\times(3\times 26)$ for the two NOT models, respectively.
For the inverse diffusion model, we generate the SDF from pure Gaussian noise on uniform, fixed grids, which is a common image-to-image translation task. The residual U-Net is a prevalent choice for such tasks and we use a similar architecture to the original paper of the diffusion model, which applies the attention mechanism at the bottom two levels of the U-Net. To incorporate the condition of the target stress-strain curve with a shape of $51\times 1$ into the noise estimator, we use an MLP to encode the target curve and then expand and repeat the output to the same shape as the noise image, so that it can be concatenated with the image at each level of the U-Net. The output of the diffusion model is the SDF with a shape of $120\times 120$, which is then fed into the marching algorithm to extract the geometry.
4.4 Training protocol
Before training, we normalize the SDF, stress, and displacement to zero mean and unit variance. The dataset is split into 80% for training and 20% for testing. The mean square error is used as the loss function. We implement this work using the PyTorch framework, and training is performed on a single NVIDIA H100 GPU with 90GB memory on the DeltaAI machine available at NCSA. We use the Adam optimizer [44] with an initial learning rate of $0.001$, leveraging its adaptive moment estimation for stable and efficient convergence. To further enhance training stability and adapt the learning rate dynamically, we incorporate the ReduceLROnPlateau scheduler, which reduces the learning rate when the validation loss stagnates. For padding query points, a mask mechanism is applied to exclude them from the loss calculation. The loss value evolution during training is shown in Fig. S4 and Fig. S5. Detailed training information is provided in the Supplementary Information. Specifically, the computational efficiency is discussed in Supplementary Section S5.
4.5 Inference protocol
All inference tasks are performed on a single NVIDIA A100 GPU on the DELTA machine at NCSA.
For the inverse diffusion model, we generate multiple geometries for a single target stress-strain curve. We apply a filter to remove the samples with “islands” and ensure that the geometry is periodic. The filtered geometry is then fed into the forward model to predict the stress-strain curve. The predicted stress-strain curve is compared with the target curve using the $L_{2}$ relative error.
4.6 Metrics
We use the $L_{2}$ relative error as the metric to evaluate the performance of the models. For the stress-strain curve, the $L_{2}$ relative error is defined as
$L_{2}=\frac{\left\|\sigma_{eff}^{true}-\sigma_{eff}^{pred}\right\|_{2}}{\left\|\sigma_{eff}^{true}\right\|_{2}},$ (10)
where $\left\|\cdot\right\|_{2}$ is the $L_{2}$ norm. For the solution fields, the $L_{2}$ relative error is defined as
$L_{2}=\frac{1}{3}\left(\frac{\left\|\sigma_{mises}^{true}-\sigma_{mises}^{pred}\right\|_{2}}{\left\|\sigma_{mises}^{true}\right\|_{2}}+\frac{\left\|u_{x}^{true}-u_{x}^{pred}\right\|_{2}}{\left\|u_{x}^{true}\right\|_{2}}+\frac{\left\|u_{y}^{true}-u_{y}^{pred}\right\|_{2}}{\left\|u_{y}^{true}\right\|_{2}}\right),$ (11)
where the $L_{2}$ relative error is calculated over all the strain steps.
5 Data availability
The dataset is available on Zenodo [45] and the trained models are available at the Github repository https://github.com/QibangLiu/SDFGeoDesign
6 Code availability
The codes for training and inference are available at the Github repository https://github.com/QibangLiu/SDFGeoDesign
7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research is a part of the Delta research computing project, which is supported by the National Science Foundation (award OCI 2005572) and the State of Illinois, as well as the Illinois Computes program supported by the University of Illinois Urbana-Champaign and the University of Illinois System.
8 Author contributions statement
Q. Liu conceived the concept, generated the dataset, developed the methodology, implemented the models, performed the analysis, and drafted, reviewed, and edited the manuscript. S. Koric contributed to the results discussion and to the configuration of Abaqus on HPC, and reviewed and edited the manuscript. D. Abueidda and H. Meidani contributed to the result discussion and reviewed and edited the manuscript. P. Geubelle proposed the concept, supervised the project, contributed to the result discussion, and reviewed and edited the manuscript.
9 Competing interests statement
The authors declare no competing financial or non-financial interests.
S0 Overview
We here provide further details regarding the data generation procedure, model implemtation, and futher results. More details can be found in the Github repository https://github.com/QibangLiu/SDFGeoDesign
S1 Binary pixel-based vs. SDF-based geometry representation
The binary pixel-based geometry representation is commonly used for metamaterial design, where the geometry is represented by a binary image with pixel values of 0 or 1. As shown in Fig. S1(a), this representation is simple and intuitive but has limitations, such as jagged boundaries and the need for additional boundary smoothing. In contrast, the signed distance function (SDF) representation, shown in Fig. S1(b), provides a smooth boundary representation that eliminates jagged edges and simplifies the finite element simulation and 3D printing process. The SDF representation is defined as the shortest distance from a point in space to the boundary of the geometry, with negative values inside the geometry, positive values outside the geometry, and zero values on the boundary. This continuous and differentiable representation is suitable for neural network-based design and simulation tasks.
Figure S1: Geometry represented by pixel density (a) and by SDF (b).
S2 Data generation
S2.1 Periodic unit cell
We generate a random periodic unit cell by sampling a periodic 2D Gaussian random field $U\left(x\right)$ on a square domain $[0,1~{}\text{mm}]^{2}$ with a resolution of $64\times 64$, using the periodic Fourier method [2]:
$U\left(x\right)=\sum_{i=1}^{N}\sqrt{2S(k_{i})\Delta k}\left(Z_{1,i}\cdot\cos\left(\left\langle k_{i},x\right\rangle\right)+Z_{2,i}\cdot\sin\left(\left\langle k_{i},x\right\rangle\right)\right),$ (S1)
where $S$ is the spectrum of the Gaussian covariance model, $Z_{1,i},Z_{2,i}\sim N(0,1)$ are mutually independent and drawn from a standard normal distribution, and $k_{i}$ is the equidistant Fourier grid. We use the open-source Python package GSTools [1] to generate such 2D Gaussian random fields. For the Gaussian covariance model, we set the variance to 10 and the length scale to 0.15. For the Fourier space, we use $32^{2}$ modes. The GRF is evaluated as follows:
x = np.linspace(0, 1, 64)
y = np.linspace(0, 1, 64)
model = gstools.Gaussian(dim=2, var=10, len_scale=0.15)
srf = gstools.SRF(
model,
generator=”Fourier”,
period=(1,1), # periodic in x and y
mode_no=32,
seed=None)
grf=srf((x, y), mesh_type=”structured”)
For the implementation of the 2D marching algorithm, we use the open-source Python package scikit-image.
For each generated unit cell, we store the coordinates of the vertices and the contour connectivities (outer boundary and holes) in a json file, which can be read by Abaqus using the Python API for building geometry and meshing. To evaluate the SDF for each unit cell, we first build a polygon object using shapely.geometry from the open-source Python package shapely, then calculate the shortest distance of $120\times 120$ uniform grid points in $[-0.1~{}\text{mm},1.1~{}\text{mm}]^{2}$ to the polygon.
S2.2 Material properties
The material adopted in this work is acetal homopolymer resin [3] and an elastic-plastic model is used to describe its constitutive behavior. The Young’s modulus of the resin is 2.3 GPa and the Poisson’s ratio is 0.35. The plastic behavior is described by a piecewise linear hardening model presented in Table S1 and the corresponding stress-strain curve is shown in Fig. S2.
Table S1: Material properties of the homopolymer resin.
Plastic strain [%] 0.0 0.1133 0.4183 0.80645 1.2557 2.0035 3.0689
3.8873 4.7114 6.083 7.4477 8.799 11.457 12.76
Stress [MPa] 40.62 45.24 52.62 58.00 61.87 65.81 69.19
71.06 72.61 74.82 76.74 78.46 81.58 83.00
Figure S2: Stress-strain curve for homopolymer resin.
S2.3 Mesh sensitivity analysis
We conducted a mesh sensitivity analysis to determine the optimal mesh size for the finite element simulations. It is important to note that the domain boundary consists of small segments, each approximately 1/64 mm in length. This segment length corresponds to the resolution of the Gaussian random field (GRF) used to generate the geometry. Consequently, the mesh size around the contour does not exceed the segment length. Our analysis indicates that a global mesh size of 0.04 mm is sufficient to capture the compression response of the unit cells. The stress-strain curves obtained using a mesh size of 0.04 mm are compared with the curves obtained using mesh sizes of 0.01 mm, 0.02 mm, and 0.08 mm as shown in Fig. S3. The results indicate that the stress-strain curves obtained using the four mesh sizes are almost identical, confirming that a mesh size of 0.04 mm is adequate to capture the material response of the unit cell.
Figure S3: Mesh sensitivity analysis for two unit cells under compression: (a) without self-contact and (b) with self-contact.
S2.4 Data augementation
Based on the random, arbitrary periodic unit cell generation and the corresponding FE simulation, we generate 10K samples. Training the forward neural networks exhibits severe overfitting due to the limited number of training samples and the complexity of the data. To alleviate this issue, we apply data augmentation to the dataset by randomly shifting the unit cell in the $x$-direction multiple times. The new unit cell $[x_{0},x_{0}+1]\times[0,1]~{}\text{mm}^{2}$ is created by shifting the original unit cell horizontally by $x_{0}$, where $x_{0}$ is randomly selected from $[0.1,0.9]$ mm. The solution stress and displacement fields of the new unit cell are the same as the original one but with the node coordinates shifted by $x_{0}$:
$x_{\text{new}}=\begin{cases}x_{\text{original}}+1-x_{0}&\text{if }x_{\text{original}}<x_{0},\\ x_{\text{original}}-x_{0}&\text{if }x_{\text{original}}\geq x_{0}.\end{cases}$ (S2)
The stress-strain curve of the new unit cell remains the same as the original one due to the periodic boundary condition. For each unit cell, we apply 8 random shifts. After filtering out unit cells with “islands”, we obtain a total of 73,879 training samples.
S3 Forward models
S3.1 Model architectures
The forward neural operator transformer (NOT) model is designed to predict the stress-strain curve and the solution fields (Mises stress and two displacements). We summarize the most relevant hyperparameters of the two forward NOT models in this section. The NOT model consists of two main components: a geometry encoder and a solution decoder. The geometry encoder encodes the geometry represented by the signed distance function (SDF) into a latent space, which is fed into attention blocks as the Key and Value. The geometry encoder is summarized in Table S2.
Table S2: Geometry encoder architecture of the forward NOT model.
Layer # Layer type (Description) Output shape
1 Input (B, 1, 120, 120)
2 Conv2D (B, 8, 120, 120)
3 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 8, 120, 120)
4 Conv2d (DownSample) (B, 8, 60, 60)
5 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 16, 60, 60)
6 Conv2d (DownSample) (B, 16, 30, 30)
7 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 32, 30, 30)
8 Conv2d (DownSample) (B, 32, 15, 15)
9 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 64, 15, 15)
10 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 64, 15, 15)
11 Self-Attention Block (B, 64, 15, 15)
12 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 64, 15, 15)
13 Concatenate (B, 128, 15, 15)
14 SiLU+ Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 64, 15, 15)
15 Conv2D (UpSample) (B,64,30,30)
16 Concatenate (B, 96, 30, 30)
17 SiLU+ Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 32, 30, 30)
18 Conv2D (UpSample) (B, 32, 60, 60)
19 Concatenate (B, 48, 60, 60)
20 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 16, 60, 60)
21 Conv2D (UpSample) (B, 16, 120, 120)
22 Concatenate (B, 24, 120, 120)
23 SiLU+Conv2d+GroupNorm+SiLU+Conv2D (Residual Conv Block) (B, 8, 120, 120)
24 Conv2D+SiLU (B, 64, 60, 60)
25 Conv2D+SiLU (B, 64, 30, 30)
26 Conv2D+SiLU (B, 64, 15, 15)
27 Reshape + Permute (Output) (B,225,64)
Table S3: Solution decoder architecture of the forward NOT model. $N$ represents the number of query points. For the stress-strain curve prediction, $N=51$, while for solution fields prediction, $N$ represents the maximum number of sampled nodes of in the batch. $C_{in}$ represents the input channels and is equal to 1 for stress-strain curve prediction and 2 for solution fields prediction. $C_{out}$ corresponds to the output channels and is equal to 1 for stress-strain curve prediction, and $3\times 26$ for solution field prediction (Mises stress and displacements at 26 strain steps).
Layer # Layer type (Description) Output shape
1 Input (B, N, $C_{in}$)
2 Nerf Positional Encoding+Linear+ReLU (B, N, $C_{in}*31$)
3 MLP with ReLU (channels: $C_{in}*31\rightarrow 192\rightarrow 256\rightarrow 128\rightarrow 64$) (B, N, 64)
4 Cross Attension (K,V:(B,225,64), Q(B,N,64) ) (B, N, 64)
5 Cross Attension (K,V:(B,225,64), Q(B,N,64) ) (B, N, 64)
6 MLP with ReLU (channels: $64\rightarrow 128\rightarrow 256\rightarrow 256$) (B, N, 256)
7 Linear (Output) (B, N, $C_{out}$)
S3.2 Loss function
The mean square error (MSE) is used as the loss function for training the forward NOT models. The loss function is defined as
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2}.$ (S3)
When training the forward NOT model for predicting the solution fields, as the number of query points is padded to the maximum number of nodes in the batch, we apply a mask mechanism to exclude the padding points from the loss calculation:
$\text{MSE}=\frac{1}{\sum_{i=1}^{N}m_{i}}\sum_{i=1}^{N}m_{i}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (S4)
where $m_{i}$ is the mask value, which is 1 if it is not a padded point, and 0 if it is. The training hyperparameters for the forward NOT models are summarized in Table S4.
Table S4: Training hyperparameters of the forward NOT models.
Hyperparameter Value
Batch size (full field) 64
Batch size (stress-strain curve) 128
Initial learning rate 1e-3
Optimizer Adam
Scheduler ReduceLROnPlateau
Scheduler patience 20
Schedule factor 0.7
Epochs (full field) 300
Epochs (stress-strain curve) 500
Training dataset 80%
Validation dataset 20%
S3.3 Model performance
The model performance of the forward NOT model for predicting the stress-strain curve is shown in Fig. S4, where (a) displays the MSE loss of the training history and (b) presents the $L_{2}$ relative error distribution of the test data. Despite the overfitting indicated in Fig. S4(a), the $L_{2}$ relative error of the test data is as low as 2.6% with a standard deviation of 2.4%, demonstrating that the model generalizes well to the test data.
Figure S4: Performance of the NOT forward model for predicting the stress-strain curve from SDF. (a) shows the MSE loss during training, and (b) presents the $L_{2}$ relative error distribution for the test data. The mean $L_{2}$ relative error is 2.6% with a standard deviation of 2.4%.
The performance of the forward NOT model for predicting the solution fields is presented in Fig. S5. Fig. S5(a) shows the MSE loss during training, while Fig. S5(b) displays the $L_{2}$ relative error distribution for the test data. The mean $L_{2}$ relative error is 10.3% with a standard deviation of 4.6%, indicating that the model generalizes well to the test data. The best and worst prediction cases from the test data are illustrated in Fig. S6 and Fig. S7, respectively. The $L_{2}$ relative error over the total 26 strain steps is 3.9% for the best case and 72.6% for the worst case.
Figure S5: Performance of the NOT forward model for predicting Mises stress and displacement from arbitrary geometries. (a) shows the MSE loss during training, and (b) presents the $L_{2}$ relative error distribution for the test data of Mises stress and displacement. The mean $L_{2}$ relative error is 10.3% with a standard deviation of 4.6%.
Figure S6: Comparison between the Mises stress and displacement field predictions with the corresponding FE ground truth at different strain $\varepsilon$ steps for the best case of the test data. The first row shows the true Mises stress under the true deformed shape, the second row shows the predicted Mises stress under the predicted deformed shape, and the third row shows the absolute error of the Mises stress under the true deformed shape. The $L_{2}$ relative error over the total 26 strain steps is 3.9%.
Figure S7: Comparison between the Mises stress and displacement field predictions with the corresponding FE ground truth at different strain $\varepsilon$ steps for the worst case of the test data. The first row shows the true Mises stress under the true deformed shape, the second row shows the predicted Mises stress under the predicted deformed shape, and the third row shows the absolute error of the Mises stress under the true deformed shape. The $L_{2}$ relative error over the total 26 strain steps is 72.6%.
S4 Inverse model
S4.1 Diffusion model theory
In this work, the classifier-free guided diffusion model [5] is trained to design the micro-structure of the unit cell of a metamaterial that can achieve a target stress-strain curve. The labeled data $y$ consists of the stress-strain curve obtained from the ABAQUS simulation of the corresponding random geometry. The corresponding SDF image serves as the design variable $\mathbf{x_{0}}$ for the diffusion model. By applying the marching algorithm on the designed SDF images to extract the contour with a value equal to 0, the boundary of the designed geometry is obtained. The model architecture is shown in Fig. 2(d). For completeness, the denoising diffusion probabilistic model [4] and the classifier-free guidance method [5] are briefly reviewed here.
Given an initial sample $\mathbf{x_{0}}$ from a prior data distribution $\mathbf{x_{0}}\sim q(\mathbf{x})$, the forward diffusion process of the denoising diffusion probabilistic model incrementally adds a small amount of Gaussian noise to the sample across $T$ steps. This creates a sequence of samples $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$. Each step in this diffusion process is regulated by a variance schedule $\left\{\beta_{t}\in(0,1)\right\}_{t=1}^{T}$,
$q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$ (S5a)
$q\left(\mathbf{x}_{1:T}|\mathbf{x}_{0}\right)=\prod_{t=1}^{T}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right).$ (S5b)
The sample $\mathbf{x}_{0}$ gradually loses its features and eventually becomes an isotropic Gaussian distribution for large $T$. This diffusion process has the elegant property that we can sample $\mathbf{x}_{t}$ at any time step $t$ using the reparametrization trick,
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (S6a)
$q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right),$ (S6b)
where $\alpha_{t}=1-\beta_{t}$,$\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$. By reversing the diffusion process, we can sample from $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and ultimately generate the true sample $\mathbf{x}_{0}$ from Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. Using Bayesian rule, the reverse posterior distribution conditioned on $\mathbf{x}_{0}$ is traced,
$q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right),\tilde{\beta}_{t}\mathbf{I}\right),$ (S7)
where $\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{t}\right)$ and $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$. In a real generative process, we cannot directly evaluate $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$ because it requires the training sample and the entire diffusion noise dataset. Instead, we propose a posterior distribution $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$ to approximate the true posterior distribution $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$,
$p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)\right).$ (S8)
The reverse denoising process is then controlled by
$p_{\theta}\left(\mathbf{x}_{0:T}\right)=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right),$ (S9)
where $\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)$ is the same as $\tilde{\beta}_{t}\mathbf{I}$, and $\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)$ has the same form as $\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)$ but with learnable parameters $\theta$,
$\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)\right).$ (S10)
The optimization objective of the denoising diffusion probabilistic model is to maximize the log-likelihood $\log p_{\theta}\left(\mathbf{x}_{0:T}\right)$, which results in the following loss function:
$L=\left|\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)-\epsilon_{t}\left(\mathbf{x}_{t},t\right)\right|.$ (S11)
Training a NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ and minimizing the loss function $L$ leads to $\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)\approx\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)$, thus making the proposed posterior distribution $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$ close to the true posterior distribution $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$. The above reverse generative process is random and not controlled by any specific target. In specific design tasks, we aim to generate fields that represent the target, such as the force-displacement curve, which requires training the NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ with conditional information. To incorporate the condition information $\mathbf{y}$ into the diffusion process, Ho and Salimans [40] proposed a classifier-free guidance method that incorporates the scores from a conditional diffusion model $p_{\theta}(\mathbf{x}|\mathbf{c})$ and an unconditional diffusion model $p_{\theta}(\mathbf{x})$. The noise estimators $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ of $p_{\theta}(\mathbf{x})$ and $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ of $p_{\theta}(\mathbf{x}|\mathbf{c})$ are trained in a single NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)$. Here, $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ is trained with paired data $(\mathbf{x},\mathbf{c}=\mathbf{y})$, and $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ is trained with data $\mathbf{x}$ only, i.e.,$(\mathbf{x},\mathbf{c}=\emptyset)$. During the training process, the condition information $\mathbf{c}$ is randomly set as $\mathbf{c}=\emptyset$ or $\mathbf{c}=\mathbf{y}$ for sample $(\mathbf{x},\mathbf{y})$ at different time steps $t$, allowing the model to learn to generate fields both conditionally and unconditionally. During the reverse inference process, $\epsilon_{\theta}$ in Eq. S10 is replaced by the linear summation of conditional and unconditional noise estimators,
$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)=(1+w)\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)-w\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\emptyset\right),$ (S12)
where $w\geq 0$ is the guidance weight.
S4.2 Model architectures
The classifier-free guided diffusion model is designed to predict the SDF of the unit cell of metamaterial that can achieve a target stress-strain curve. The model architecture is summarized in Table S5. It consists of three main components: a stress-strain (SS) curve encoder, a time encoder, and a U-Net encoder-decoder. The outputs of the SS curve encoder and the time encoder are fed into the residual blocks in the U-Net.
Table S5: Geometry encoder architecture of the forward NOT model (B is batch size).
Network Layer # Layer type (Description) Output shape
Encoding 1 Input (SS curve) (B, 51)
stress-strain 2 MLP with SiLU (B, 32)
(channels: 51323232absent323232$\rightarrow$ 32$\rightarrow$ 32$\rightarrow$ 32→ 32 → 32 → 32)
3 Linear (Output) (B, 16)
Encoding 1 Input (time t) (B, )
time t 2 Time embedding+Linear (Output) (B, 64)
U-Net 1 Input (SDF at time t) (B, 1,120,120)
encoder 2 Conv2D (B, 16, 120, 120)
3 Residual Conv Block (B, 16, 120, 120)
4 Conv2d (DownSample) (B, 16, 60, 60)
5 Residual Conv Block (B, 32, 60, 60)
6 Conv2d (DownSample) (B, 32, 30, 30)
7 Residual Conv Block (B, 64, 30, 30)
8 Conv2d (DownSample) (B, 64, 15, 15)
9 Residual Conv Block + self-attention (B, 128, 15, 15)
U-Net 10 Residual Conv Block (B, 128, 15, 15)
bottle-neck + self-attention + Residual Conv Block
U-Net 11 Concatenate (B, 256, 15, 15)
decoder 12 Residual Conv Block + self-attention (B, 128, 15, 15)
13 Conv2D (UpSample) (B, 128, 30, 30)
14 Concatenate (B, 192, 30, 30)
15 Residual Conv Block + self-attention (B, 64, 30, 30)
16 Conv2D (UpSample) (B, 64, 60, 60)
17 Concatenate (B, 96, 60, 60)
18 Residual Conv Block (B, 32, 60, 60)
19 Conv2D (UpSample) (B, 32, 120, 120)
20 Concatenate (B, 48, 120, 120)
21 Residual Conv Block (B, 16, 120, 120)
22 GroupNorm+SiLU+Conv2D (Output) (B, 1, 120, 120)
1 Inputs
Residual (image, encoded SS curve, encoded t) (B, in, W, H), (B,16), (B,64)
Conv Block 2 Conv2D,SiLU+Linear, SiLU+Linear (B, out, W, H), (B,out), (B,out)
3 –,Expand+repeat,Expand+repeat (B, out, W, H)×$\times 3$
4 Concatenate (B, 3*out, W, H)
5 Conv2d (B, out, W, H)
6 Residual sumation (B, out, W,H)
The overview of the training hyperparameters of the classifier-free guided diffusion model is shown in Table S6.
Table S6: Training hyperparameters of the classifier-free guided diffusion model.
Hyperparameter Value
Batch size 64
Initial learning rate 1e-3
Optimizer Adam
Scheduler ReduceLROnPlateau
Scheduler patience 20
Schedule factor 0.7
Epochs 300
Training dataset 80%
Testing dataset 20%
Time steps ($T$) 500
S4.3 Model performance
The model performance of the classifier-free guided diffusion model is shown in Fig. S8
Figure S8: MSE loss of training history of the diffusion model.
S5 Computational efficiency
The FE simulations were executed on CPUs of the Delta machine at the National Center for Supercomputing Applications (NCSA) at the University of Illinois at Urbana-Champaign. Model training was conducted on a single NVIDIA H100 GPU on the DeltaAI machine at NCSA. Inference tasks for the forward NOT model and the inverse design model were performed on an NVIDIA A100 GPU on the Delta machine. The computational efficiency of data generation, model training, and inference is summarized in Table S7.
Table S7: Computational efficiency of data generation, model training, and inference.
Task Time
FE simulation (1 cpu) 2.3 min/sample
NOT training (SS curve) 16 s/epoch
NOT training (full field) 96 s/epoch
Diffusion model training 47 s/epoch
NOT inference (SS curve) 1.6e-4 s/solution
NOT inference (full field) 1.0e-2 s/solution
Inverse design 0.46 s/solution